OpenStudy (anonymous):

Find the magnitude and direction angle θ of the following vectors. Round the magnitude to the nearest tenth, and round the direction angle to the nearest degree, if rounding is necessary. u = (9, -2) Choice 'A' |u| ≈ 9.2; θ ≈ 12° Choice 'B' |u| ≈ 12.2; θ ≈ -13° Choice 'C' |u| ≈ 9.2; θ ≈ -13° Choice 'D' |u| ≈ 13.2; θ ≈ 24° @jdoe0001 @whpalmer4 @Loser66 @agent0smith

OpenStudy (johnweldon1993):

\[\large \text{Magnitude} = \sqrt{x^2 + y^2}\] \[\large \text{Direction} = \theta = \tan^{-1}\frac{y}{x}\]